# FluidNexus: 3D Fluid Reconstruction and Prediction from a Single Video

### v1

@@ 1 Introduction
We propose a novel video-to-prediction framework FluidNexus to address these challenges. Our system consists of two components: (1) A novel-view video synthesizer, comprising a frame-wise novel-view diffusion model and a video diffusion refiner, which generates multi-view videos from a single-view input. The frame-wise model synthesizes individual frames from novel viewpoints without accounting for fluid dynamics, while the video diffusion refiner enhances these frames into a coherent and realistic video sequence. (2) A two-layer particle representation, which bridges differentiable physics and differentiable rendering using two sets of moving particles, reconstructs 4D fluid flow motion from multi-view video input. Specifically, our two-layer particle model integrates a group of physical particles, which implements a differentiable physics simulator based on Position-Based Fluid (PBF) [37], with a group of visual particles, which establishes a differentiable rendering pipeline leveraging 3D Gaussian Splatting [28] to link the input videos with the 4D reconstructions. The interaction between these two particle groups enables a sparse yet effective representation of the spatiotemporal fluid flow during reconstruction based on multi-view video input, while simultaneously enforcing physical constraints on fluid motions during prediction, where ground truth video reference is no longer available.
@@ 3 Approach
Figure 2: FluidNexus in reconstruction. From a single video, we synthesize multiple novel-view videos as references for 3D fluid reconstruction. We then sequentially optimize the two-layer particle fluid representations over time, using the multi-view video frames to compute the visual loss and the physics constraints to compute the physics loss. Our reconstruction output is the 3D fluid appearance and velocity fields over all input frames.
@@
We propose FluidNexus, which consists of a novel-view video synthesizer (left of Figure 2) to create temporally consistent multi-view videos based on a single video input and a physics-integrated two-layer particle representation (right of Figure 2) to reconstruct and predict fluid motion based on multi-view video inputs. We elaborate these two components in the following.
@@ 3.1 Novel-view Video Synthesizer
3.2 Two-layer Particle Fluid Representation
@@
Our fluid representation consists of two types of particles: physical particles that represent the fluid velocity field and density field in ambient 3D space, and visual particles that represent the fluid appearance and are passively advected by the velocity field. Physical particles at any timestep $t$ are defined by their positions $\mathbf{p}_{t}\in\mathbb{R}^{3N_{t}^{\text{physical}}}$ where $N_{t}^{\text{physical}}$ denotes the physical particle count at $t$, and the associated particle velocities $\mathbf{u}_{t}\in\mathbb{R}^{3N_{t}^{\text{physical}}}$ (not to be confused with the velocity field $\mathbf{V}_{t}:\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$). The velocity at any 3D point $\mathbf{x}$ at timestep $t$ is computed through kernel-weighted interpolation:
@@ Appendix B Technical Details
Figure S1: Flowchart of FluidNexus. The reconstruction algorithm is shown in Alg. S1. On the left is the reconstruction process which takes a single video as input and produces the two-layer particles within the video duration. On the right is the prediction process which takes the reconstructed two-layer particles as input and produces the particles in future timesteps.

### v2

@@ 1 Introduction
We propose a novel video-to-prediction framework FluidNexus to address these challenges. Our system consists of two components: (1) A novel-view video synthesizer, comprising a frame-wise novel-view diffusion model and a video diffusion refiner, which generates multi-view videos from a single-view input. The frame-wise model synthesizes individual frames from novel viewpoints without accounting for fluid dynamics, while the video diffusion refiner enhances these frames into a coherent and realistic video sequence. (2) A visual-physical particle representation, which bridges differentiable physics and differentiable rendering using two sets of moving particles, reconstructs 4D fluid flow motion from multi-view video input. Specifically, our visual-physical particle model integrates a group of physical particles, which implements a differentiable physics simulator based on Position-Based Fluid (PBF) [37], with a group of visual particles, which establishes a differentiable rendering pipeline leveraging 3D Gaussian Splatting [28] to link the input videos with the 4D reconstructions. The interaction between these two particle groups enables a sparse yet effective representation of the spatiotemporal fluid flow during reconstruction based on multi-view video input, while simultaneously enforcing physical constraints on fluid motions during prediction, where ground truth video reference is no longer available.
@@ 3 Approach
Figure 2: FluidNexus in reconstruction. From a single video, we synthesize multiple novel-view videos as references for 3D fluid reconstruction. We then sequentially optimize the visual-physical particle fluid representations over time, using the multi-view video frames to compute the visual loss and the physics constraints to compute the physics loss. Our reconstruction output is the 3D fluid appearance and velocity fields over all input frames.
@@
We propose FluidNexus, which consists of a novel-view video synthesizer (left of Figure 2) to create temporally consistent multi-view videos based on a single video input and a physics-integrated particle representation (right of Figure 2) to reconstruct and predict fluid motion based on multi-view video inputs. We elaborate these two components in the following.
@@ 3.1 Novel-view Video Synthesizer
3.2 Visual-Physical Particle Fluid Representation
@@
Our fluid representation consists of two types of particles: physical particles that represent the fluid velocity field and density field in ambient 3D space, and visual particles that represent the fluid appearance and are passively advected by the velocity field. We need two different types of particles because intrinsic difference between the distributions: the velocity is defined in ambient 3D space while the appearance (visible fluid) is generally defined only in specific regions.
Physical particles at any timestep $t$ are defined by their positions $\mathbf{p}_{t}\in\mathbb{R}^{3N_{t}^{\text{physical}}}$ where $N_{t}^{\text{physical}}$ denotes the physical particle count at $t$, and the associated particle velocities $\mathbf{u}_{t}\in\mathbb{R}^{3N_{t}^{\text{physical}}}$ (not to be confused with the velocity field $\mathbf{V}_{t}:\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$). The velocity at any 3D point $\mathbf{x}$ at timestep $t$ is computed through kernel-weighted interpolation:
@@ Appendix B Technical Details
Figure S1: Flowchart of FluidNexus. The reconstruction algorithm is shown in Alg. S1. On the left is the reconstruction process which takes a single video as input and produces the visual-physical particles within the video duration. On the right is the prediction process which takes the reconstructed visual-physical particles as input and produces the particles in future timesteps.